# Reassessing Collaborative Writing Theories and Frameworks in the Age of LLMs: What Still Applies and What We Must Leave Behind

### v2

@@ Abstract.
In this paper, we conduct a critical review of existing theories and frameworks on human-human collaborative writing to assess their relevance to the current human-AI paradigm in professional contexts, and draw seven insights along with design implications for human-AI collaborative writing tools. We found that, as LLMs nudge the writing process more towards an empirical “trial and error” process analogous to prototyping, the non-linear cognitive process of writing will stay the same, but more rigor will be required for revision methodologies. This shift would shed further light on the importance of coherence support, but the large language model (LLM)’s unprecedented semantic capabilities can bring novel approaches to this ongoing challenge. We argue that teamwork-related factors such as group awareness, consensus building and authorship – which have been central in human-human collaborative writing studies – should not apply to the human-AI paradigm due to excessive anthropomorphism. With the LLM’s text generation capabilities becoming essentially indistinguishable from human-written ones, we are entering an era where, for the first time in the history of computing, we are engaging in collaborative writing with AIs at workplaces on a daily basis. We aim to bring theoretical grounding and practical design guidance to the interaction designs of human-AI collaborative writing, with the goal of enhancing future human-AI writing software.
@@ 1. Introduction
As large language models (LLMs) increasingly become our daily co-writers in a uniquely asymmetrical relationship unlike any we have previously encountered, there is a need to reassess foundational theories and frameworks of collaborative writing to adapt to the human-AI paradigm. Their capabilities in natural language generation are becoming indistinguishable from human writing (Hämäläinen et al., 2023; Clark et al., 2021), but the mechanism of human-AI collaborative writing is clearly distinct from a human-human one. To reflect this paradigm shift, new frameworks and taxonomies are emerging, from domain-specific taxonomy development (Lee et al., 2024b), recognizing human-side effort through entropy and information gain (Wan et al., 2024), getting inspiration from the Overall Equipment Effectiveness Model (Wilbers et al., 2024), to exploring a non-linear framework to go beyond the linear human-only/human-AI/AI-only dimension of collaboration modes. (Hutson, 2025). While many of these proposals build on a partial aspect of existing theories and frameworks of collaborative writing, we believe that this must be approached holistically, synthesizing what past studies have taught us about the mechanisms of collaborative writing. It goes without saying that legacy theories and frameworks on collaborative writing have (naturally) focused only on the human-human mode. While the already established understandings of human-human collaborative writing should certainly apply to human-AI to some degree, we believe now is the moment to reassess our understanding of collaborative writing so far, and adapt it to the ever increasing human-AI mode.
@@
In this paper we conduct a critical review of legacy theories and frameworks on collaborative writing, and analyze what we see to be relevant to the human-AI paradigm as well as what only applies to the human-human paradigm, bringing clarity to the theoretical groundings required for the design of future human-AI collaborative writing tools. Our main findings are that, given the shift in the human’s role from hands-on writer to an editorial decision maker, the non-linear cognitive process of writing will stay the same, but more attention will be required for rigorous revision methodologies. In turn, this may reduce the focus on the planning aspects of collaborative writing, nudging writing more towards an empirical “trial and error” process than a fully planned out one, analogous to prototyping. This shift would require more focus on supporting coherence between human and AI writing, but we believe that the LLM’s unprecedented semantic capabilities can bring a break through to this ongoing challenge. Teamwork-related factors such as group awareness, consensus building and authorship, which have been central in human-human collaborative writing studies, will not apply to the human-AI paradigm due to excessive anthropomorphism, but it is equally the AI co-writer’s role to facilitate those factors between human writers to offload any unnecessary cognitive load. Finally, be it human-written or AI-written, writing will remain to be socially constructed by readers, which suggests that the consideration of reader profiles is inseparable from the text generation process. We provide practical design implications and example interface sketches with our insights to provide concrete design directions.
@@ 1.1. Study scope
The scope of this paper is limited to the pragmatic support of writing tasks in organizational workplace settings. This includes cases such as clinicians drafting discharge summaries, course coordinators drafting a response to student questions on online forums, meeting participants drafting a summary to send out to their team, or academic staff crafting grant proposals. We do not consider in fictional writing or rhetoric, nor do we consider topics around authorship, emotional satisfaction, or the deskilling aspects of writing. In particular, collaborative writing has traditionally had much focus on ownership and authorship, but building on Schneiderman (Shneiderman, 2022) we take the stance that LLMs are and should merely be a tool to help us get the job done in the most effective manner, and avoid any unnecessary forms of anthropomorphism in framing the AI as a “team mate”.
@@ 2.1. Social Computing Paradigm
The Computers As Social Actors (CASA) paradigm (Nass et al., 1994), which revealed that humans apply the same social rules they would apply to humans to computers, has been a dominant framework in Human Computer Interaction (HCI). The underlining premise is simple: the way humans interact with computers is inherently social. The social response theory (Nass and Moon, 2000) took this a step further, explaining that such social rules are “mindlessly applied”, and that alternative explanations such as anthropomorphism simply cannot explain this phenomenon. Hence, to design interactions for new technologies, it is best to reframe computer interaction as a social interaction, and appropriately lean on relevant studies in social sciences along the way.
@@
Recently, such an approach has begun to become evident in dialogue studies, as represented in the emerging field of Human-Machine Communication (HMC), an area of communication research that regards technology as “communicative subjects, instead of mere interactive objects” (Guzman and Lewis, 2020). That is, HMC treats machines no longer as mediators, but as actors directly exchanging messages with humans. This reframing expands the focus of communication studies from human-human only to include human-AI contexts. For example, while it is well known that the lexical, syntactic and semantic alignment between two people engaged in a dialogue is crucial for mutual understanding (Pickering and Garrod, 2004), Ostrand and Berger claim that such alignment should be properly considered when building LLMs as well (Ostrand and Berger, 2024) for improved usability and personalization. This has also led to interesting discussions on the effects such alignment would have on bias amplification and echo chambers (Knoeferle et al., 2025).
With collaborative writing, however, there is still little work at the intersection of traditional human-human collaborative writing and the emerging human-AI writing paradigm. This is perhaps due to the fact that, until recently, writing with AI primarily referred to assistance tools such as spell- or grammar-checking. It also remains unclear as to what extent writing with LLMs can be considered “collaborative writing”. However, when a human tasks an LLM to write up a draft of a document, the human-AI pair is working “toward coordinated consensus that is reflected in a document that is written by one of the team members”, which we see to be undeniably qualifying as a form of collaborative writing (Lowry et al., 2004) (it is also worth noting that, according to Posner and Baecker (Posner and Baecker, 1992), single writer strategy was actually the most common strategy adopted even in human-only groups). If that is the case, the underlining reframing that is needed here remains the same. Writing is hardly a solitary act, be it with humans or with AI: it happens through the participation and interaction with others (Ede and Lunsford, 2006). We can see here that writing, too, is inherently social. As Wang et al. (Wang et al., 2020) suggest, it is crucial to integrate existing theories from HCI and Computer-Supported Cooperative Work (CSCW) in designing for the optimal form of human-AI collaborative writing.
@@ 2.2. Collaborative Writing
In 1969, Michel Foucault questioned the notion of the solitary author in the book “What is an author?”, stating that the author’s role is to regulate meaning and the circulation of text in an ever changing society (Foucault, 1979). The cognitive process theory of writing (Flower and Hayes, 1981), a widely adopted theory for unveiling an individual’s writing process, was updated by Hayes 15 years later (Hayes, 1996) to add more emphasis to the “social” aspect of writing, incorporating readers and collaborators into their framework. Ede and Lunsford (Lunsford and Ede, 1986) were the among the first to initiate collaborative writing as an academic research topic in 1986, reporting that 87 percent of respondents from six major professions sometimes wrote as a team or a group. Through further analysis they found that there were two distinct modes in collaborative writing: the “hierarchical” mode where one person has control and requires structure and division of labor, and the “dialogic” mode where all participants have equal input with shared authorship, which requires consensus and compromise (Ede and Lunsford, 2006). However, they also quickly realized that there was a significant lack of shared consensus on what collaborative writing is. As Allen et al. (Allen et al., 1987) explain: “very little detail is known about collaborative writing processes in general”. Due to the highly complex nature of document structures as well as the strong connection between writing and subjective emotions, collaborative writing is perhaps one of the most difficult forms of collaboration. As Farkas explains, there may not even be a common language for collaborative writing beyond a certain level (Lay and Karis, 1991). In light of this, Lowry et al. (Lowry et al., 2004) established the taxonomy of collaborative writing, outlining not just the various roles but also the writing strategies that groups adopt when writing together. As the role of computers and software becomes increasingly important in writing, this taxonomy is still getting updated to this day, including Collage (Buschek, 2024), CoCo matrix (Wan et al., 2024), Overall Writing Effectiveness (Wilbers et al., 2024), neurodiversity-inspired non-linear framework (Hutson, 2025), taxonomy of “text functions” (Larsen-Ledet et al., 2020) to name a few. Our work builds on these studies of emerging taxonomies, clarifying the limitations of existing human-human taxonomies and suggesting what needs to be renewed.
LLMs have brought a turning point to this field with their impressive text generation and semantic analysis capabilities. Given that writing tasks in educational and communication contexts are the second most frequent usage of Claude (Handa et al., 2025), an unprecedented number of workers are engaging in some form of collaborative writing with AIs on a daily basis. We believe it is more important now than ever to unveil the mechanisms of collaborative writing, especially between humans and AI.
@@ 2.3. Human-AI collaboration
Achieving synergy between human and AI performance has been the core focus of human-centered AI and human-robot interaction studies. One example is Explainable AI (XAI), attempting to achieve complementary performance between humans and AI for various tasks such as decision making and classification, through the transparency of the otherwise black box AI. Allowing users to drill down and ask follow up questions until they are satisfied has been an often discussed yet technically difficult challenge in XAI (Weld and Bansal, 2019), but LLMs have dramatically enabled us to study this approach further.
@@
In general, however, the various approaches in XAI have led to mixed results with regards to its effectiveness on performance (Miller, 2023), which leads to Vaccaro et al.’s (Vaccaro et al., 2024) impressive meta-analysis suggesting that creation-related tasks (including writing) with GenAI statistically have more potential for achieving synergy– the positive complementary outcome from human-AI collaboration – and that the appropriate allocation of tasks between humans and AI may be the key to further progression. This is in line with past work (Zhang et al., 2020) which found that showing AI confidence or local feature explanations had no or detrimental effect to decision accuracy, suggesting that XAI can be most effective when the error boundaries of human’s and AI’s are as distinct as possible.
@@
In this sense, we believe that the field of collaborative writing has much room for development when viewed from the human-AI collaboration lens. AI-assisted features such as auto-complete and the real-time collaboration capabilities of tools such as Google Docs or Overleaf have been studied broadly, but when it comes to collaborating with AI as a co-writer, we believe the appropriate understanding of their uniqueness and how they tie in to the already established understandings of human-human collaborative writing is yet to be defined. Similar to Yang et al. (Yang et al., 2022), we intentionally separate AI-assisted writing from AI as the co-writer.
@@
Existing studies for AI as a co-writer have mostly focused on creative writing (Clark et al., 2018; Yang et al., 2019; Calderwood et al., 2020; Yang et al., 2022). Clark et al. (Clark et al., 2018) suggests that by reducing the intrusiveness of AI assistance for creative writing and adopting a pull method for the interactions, it can “allow people to write more closely to their normal writing process.” Yang et al. (Yang et al., 2022) took a turn-taking approach in building a “finish the story” system, finding that when users are equipped with the ability to edit AI generated text, they require only partial readability for the AI outputs because they can go in and fix the fluency of the overall text themselves. At the same time, in the adjacent field of image generation, Wadinambiarachchi et al. (Wadinambiarachchi et al., 2024) studied the effects that Midjourney111See https://www.midjourney.com/ – a generative AI system for image creation – has on non-designers and found that, without appropriate literacy on how to use the tool, GenAI support could lead to higher fixation of ideas, fewer ideas, less variety, and lower originality. This is a common challenge that is similarly seen in the field of Explainable AI, known as over-reliance (Vasconcelos et al., 2023). Due to the context-dependent nature of LLMs and GenAI, it is difficult to holistically understand the capabilities of LLM to apply to interaction designs. To tackle this, Lee et al. (Lee et al., 2022) proposes CoAuthor, a large dataset of recorded interactions between writers and GPT-3 across 1445 writing sessions, which has paved the way for HCI researchers to analyze human-AI writing cases in detail.
@@
Our review builds on this fundamental Human-AI collaboration research to draw out the best of human and AI performance, specifically for writing tasks in organizational workplace settings, and in a way that is rooted in existing theories and frameworks. To the best of our knowledge, this is the first work to holistically reassess legacy collaborative writing theories and frameworks for the LLM era.

### v3

@@ Abstract.
In this paper, we conduct a critical review of existing theories and frameworks on human-human collaborative writing to assess their relevance to the current human-AI paradigm in organizational workplace settings, and draw seven insights along with design implications for human-AI collaborative writing tools. Our main finding was that, as we delegate more writing to AI, our cognitive process shifts from the traditional planning/translating/reviewing process to a planning/waiting/reviewing process, breaking the process due to the waiting that occurs in between. To ensure that our cognitive process remains intact, we suggest a “prototyping” approach, where the tool allows for faster iterations of the cognitive process by starting with smaller chunks of text, and gradually moving on to a fully fleshed-out document. We aim to bring theoretical grounding and practical design guidance to the interaction designs of human-AI collaborative writing, with the goal of enhancing future human-AI writing software.
@@ 1. Introduction
When humans made the shift from writing on paper to typing on word-processors in the 1980s, studies found that writers engaged in less conceptual planning, reread texts more extensively, and there was poorer recall of their own text (Haas, 2013). How then, will writing with Large Language Models (LLMs) affect the way that we write? As LLMs increasingly become our daily co-writers, and as their capabilities in natural language generation become increasingly indistinguishable from human writing (Hämäläinen et al., 2023; Clark et al., 2021; Zhu et al., 2024), we see a need to reassess foundational theories and frameworks on collaborative writing, and see how they adapt to the human-AI paradigm. To reflect this paradigm shift, new frameworks and taxonomies of collaborative writing are emerging, from domain-specific taxonomy development (Lee et al., 2024b), recognizing human-side effort through entropy and information gain (Wan et al., 2024), getting inspiration from the Overall Equipment Effectiveness Model (Wilbers et al., 2024), to exploring a non-linear framework to go beyond the linear human-only/human-AI/AI-only dimension of collaboration modes. (Hutson, 2025). While many of these proposals build on a partial aspect of existing theories and frameworks on collaborative writing, we believe that this must be approached holistically, synthesizing what past studies have taught us about the mechanisms of collaborative writing. Naturally, legacy theories and frameworks on collaborative writing have all focused on humans writing with humans. To what degree do these understandings of human-human collaborative writing apply to the human-AI paradigm? We believe now is the moment to look back to the diverse literature in collaborative writing, and see what we can take from them to the increasing human-AI mode.
@@
In this paper we conduct a critical review of legacy theories and frameworks on collaborative writing, and analyze what we see to be relevant to the human-AI paradigm as well as what only applies to the human-human paradigm, bringing clarity to the theoretical groundings required for the design of future human-AI collaborative writing tools. Our main findings are that, when the human’s role shifts from a hands-on writer to an editorial decision maker, the cognitive process of writing (Flower and Hayes, 1981) shifts from the traditional planning/translating/reviewing process to planning/waiting/reviewing (Figure 1), making the cognitive process discontinuous. For this we suggest AI tools to streamline the cognitive process by shortening the waiting stage as well as accelerating the overall cycle through a “prototyping” approach to writing, such as enabling human writers to revise AI generated text faster, and directly connecting that as the next set of prompts to test new ideas. Coherence support between human and AI writing is another ongoing challenge, but the LLM’s strong semantic capabilities can provide real-time analyses that were previously difficult with traditional methods. Teamwork-related factors such as group awareness, consensus building and authorship, which have been central in human-human collaborative writing studies, will not apply to the human-AI paradigm due to excessive anthropomorphism, but it is the AI co-writer’s role to facilitate such factors for human writers to offload cognitive load. Finally, be it human-written or AI-written, writing will remain to be socially constructed by readers, which suggests that the consideration of reader profiles is inseparable from the text generation process. We provide practical design implications and example interface sketches with our insights to provide concrete design directions.
The main contributions of this work are as follows:
We define the limitations of existing human-human collaborative writing theories and frameworks when applied to the human-AI paradigm
We propose an updated cognitive model when collaborative writing with AI and discuss its implications
We provide actionable design implications for future human-AI writing tools based on what we can apply from collaborative writing theories
@@ 1.1. Study scope
The scope of this paper is limited to the pragmatic support of writing tasks in organizational workplace settings. A prototypical example of such case would be drafting emails with the help of LLMs. Other cases include clinicians drafting discharge summaries, course coordinators drafting a response to student questions on online forums, meeting participants drafting a summary to send out to their team, or academic staff crafting grant proposals. We do not consider fictional writing or rhetoric (Yang et al., 2022; Calderwood et al., 2020), nor do we consider the deskilling aspects of writing (Kosmyna et al., 2025). In particular, studies on collaborative writing have traditionally focused much on ownership and authorship(Ede and Lunsford, 2001; Lunsford and Ede, 1986; Stokel-Walker, 2023), but building on the views of Naikar et al. (Naikar et al., 2025) and Schneiderman (Shneiderman, 2022) we take the stance that LLMs are and should merely be a tool to help us get the job done in the most effective manner, and avoid any unnecessary forms of anthropomorphism that frames the AI as a “team mate”.
@@ 2.1. Social Computing Paradigm
The Computers As Social Actors (CASA) paradigm (Nass et al., 1994), which revealed that humans often apply social rules to computers, has been a dominant framework in Human Computer Interaction (HCI). The underlining premise is simple: the way humans interact with computers is inherently social. The social response theory (Nass and Moon, 2000) took this a step further, explaining that such social rules are often “mindlessly applied” (Nass and Moon, 2000), and that alternative concepts such as anthropomorphism are not sufficient to explain this phenomenon. Hence, to design interactions for new technologies, it is best to reframe computer interaction as a social interaction, and appropriately lean on relevant studies in social sciences along the way.
@@
Recently, such an approach has become evident in dialogue studies, as represented in the emerging field of Human-Machine Communication (HMC), an area of communication research that regards technology as “communicative subjects, instead of mere interactive objects” (Guzman and Lewis, 2020). That is, HMC treats machines no longer as mediators, but as actors directly exchanging messages with humans. This reframing expands the focus of communication studies beyond human-human contexts to human-AI contexts. For example, while it is well known that the lexical, syntactic and semantic alignment between two people engaged in a dialogue is crucial for mutual understanding (Pickering and Garrod, 2004), Ostrand and Berger claim that such alignment should be considered when building LLMs as well (Ostrand and Berger, 2024), for improved usability and personalization. This has also led to interesting discussions on the effects that such alignment would have on bias amplification and echo chambers (Knoeferle et al., 2025).
With collaborative writing, however, there is still little work that applies what we know from human-human collaborative writing to the emerging human-AI writing paradigm. This is partially due to the fact that, until recently, writing with AI primarily referred to limited assistance tools such as spell- or grammar-checking. LLMs have brought a turning point to this field with their impressive text generation and semantic analysis capabilities. Writing tasks in educational and communication contexts are now the second most frequent usage of Claude (Handa et al., 2025) (the first being coding). Inspired by the CASA paradigm, the way humans write with AI could also be inherently social, which leads us to this systematic review to see which aspects of existing studies on human-human writing will apply to human-AI writing.
@@ 2.2. Collaborative Writing
While writing was once thought to be a solitary act (Foucault, 1979), there was a gradual realization that, in many cases, humans in fact write together, leading to the field of collaborative writing. To reflect this, Flower and Hayes’ widely adopted theory of an individual’s cognitive writing process (Flower and Hayes, 1981) was updated by Hayes 15 years later (Hayes, 1996), emphasizing to the “social” aspect of writing and incorporating collaborators as well as readers into their framework. Ede and Lunsford (Lunsford and Ede, 1986) were the among the first to initiate collaborative writing as an academic research topic in 1986, reporting that 87 percent of respondents from six major professions sometimes wrote as a team or a group. Through further analysis they found that there were two distinct modes in collaborative writing: the “hierarchical” mode where one person has control and requires structure and division of labor, and the “dialogic” mode where all participants have equal input with shared authorship, which requires consensus and compromise (Ede and Lunsford, 2006). However, they also quickly realized that there was a significant lack of shared consensus on what collaborative writing is. As Allen et al. (Allen et al., 1987) explain: “very little detail is known about collaborative writing processes in general”. Due to the highly complex nature of document structures as well as the strong connection between writing and subjective emotions, collaborative writing is perhaps one of the most difficult forms of collaboration. As Farkas explains, there may not even be a common language for collaborative writing beyond a certain level (Lay and Karis, 1991). In light of this, Lowry et al. (Lowry et al., 2004) established the taxonomy of collaborative writing, outlining not just the various roles but also the writing strategies that groups adopt when writing together. As the role of computers and software are becoming increasingly important in writing, this taxonomy is getting updated to this day, including Collage (Buschek, 2024), CoCo matrix (Wan et al., 2024), Overall Writing Effectiveness (Wilbers et al., 2024), neurodiversity-inspired non-linear framework (Hutson, 2025), taxonomy of “text functions” (Larsen-Ledet et al., 2020) to name a few. Our work builds on these studies of emerging taxonomies, clarifying the limitations of existing human-human taxonomies and suggesting what needs to be renewed.
It also remains unclear as to what extent writing with LLMs can be considered “collaborative writing”. However, when a human tasks an LLM to write up a draft of a document, the human-AI pair is working “toward coordinated consensus that is reflected in a document that is written by one of the team members” (Lowry et al., 2004), which qualifies as a form of collaborative writing (it is also worth noting that, according to Posner and Baecker (Posner and Baecker, 1992), single writer strategy was actually the most common strategy adopted, even in human-only groups). If that is the case, it seems viable to consider writing with AI to be a legitimate form of collaborative writing. Be it with humans or with AI, writing is hardly a solitary act: it happens through the participation and interaction with others (Ede and Lunsford, 2006).
@@ 2.3. Human-AI collaboration
Achieving synergy between human and AI performance has been a primary focus of human-centered AI and human-robot interaction studies. One example is Explainable AI (XAI), which, among other goals, seeks to achieve complementary performance between humans and AI in tasks such as decision making and classification by increasing the transparency of otherwise black-box models. Allowing users to drill down and ask follow up questions until they are satisfied has been an often discussed yet technically difficult challenge in XAI (Weld and Bansal, 2019), but LLMs have dramatically enabled us to study this approach further (Ma et al., 2024).
@@
In general, however, the various approaches in XAI have led to mixed results with regards to its effectiveness on performance (Miller, 2023), which leads to Vaccaro et al.’s (Vaccaro et al., 2024) impressive meta-analysis suggesting that creation-related tasks with GenAI (including writing) have more statistical potential for achieving synergy – the positive complementary outcome from human-AI collaboration – and that the appropriate allocation of tasks between humans and AI may be the key to further progression. This is in line with past work (Zhang et al., 2020) which found that showing AI confidence or local feature explanations had no or detrimental effect to decision accuracy, suggesting that XAI can be most effective when the error boundaries of human’s and AI’s are as distinct as possible.
@@
In this sense, we believe that the field of collaborative writing has much room for development when viewed from the human-AI collaboration lens. AI-assisted features such as auto-complete and the real-time collaboration capabilities of tools such as Google Docs or Overleaf have been studied broadly, but when it comes to collaborating with AI as a co-writer, we believe the appropriate understanding of their uniqueness and how they tie in to the already established understandings of human-human collaborative writing is yet to be defined. Similar to Yang et al. (Yang et al., 2022), we intentionally distinguish AI-assisted writing from AI as the co-writer.
@@
Existing studies for AI as a co-writer have mostly focused on creative writing (Clark et al., 2018; Yang et al., 2019; Calderwood et al., 2020; Yang et al., 2022). Clark et al. (Clark et al., 2018) suggest that by reducing the intrusiveness of AI assistance for creative writing and adopting a pull method for interactions, it can “allow people to write more closely to their normal writing process.” Yang et al. (Yang et al., 2022) took a turn-taking approach in building a “finish the story” system, finding that when users are equipped with the ability to edit AI generated text, they require only partial readability from AI because they can go in and fix the fluency themselves. At the same time, in the adjacent field of image generation, Wadinambiarachchi et al. (Wadinambiarachchi et al., 2024) studied the effects that Midjourney111See https://www.midjourney.com/ – a generative AI system for image creation – has on non-designers and found that, without appropriate literacy on how to use the tool, GenAI support could lead to higher fixation of ideas, fewer ideas, less variety, and lower originality. This is a common challenge that is similarly seen in the field of Explainable AI, known as over-reliance (Vasconcelos et al., 2023). Due to the context-dependent nature of LLMs and GenAI, it is difficult to holistically understand the capabilities of LLM to apply to interaction designs. To tackle this, Lee et al. (Lee et al., 2022) proposes CoAuthor, a large dataset of recorded interactions between writers and GPT-3 across 1445 writing sessions, which has paved the way for HCI researchers to analyze human-AI writing cases in detail.
@@
Our review builds on these fundamental Human-AI collaboration research, and seek to understand how human-AI writing in organizational workplace settings can be designed to draw out the best of human and AI performance. To the best of our knowledge, this is the first work to holistically reassess legacy collaborative writing theories and frameworks for the LLM era.